# Sparse Autoencoders for Sequential Recommendation Models: Interpretation and Flexible Control

### v1

@@ 7.2. Case Studies
At first, we examine the top features found in section 7.1.1. Figure 5 presents histograms of positive per-item feature activations for features exhibiting the highest correlation with Sci-Fi and Action. Histogram colors indicate whether the activation happened on an item of the associated genre. The proportion of such items increases notably with higher activation levels, which is reminiscent of the behavior of interpretable SAE features observed in NLP models (Cunningham et al., 2023; Bricken et al., 2023; Templeton et al., 2024). We found the same pattern for all other genres. Activation of a given feature above some threshold typically indicates the presence of respective genre. However, the threshold varies between features, as well as the scale of activation values.
@@
While these features correspond to the genres as labeled in the dataset, there are features that go beyond the predefined classes.
@@
To investigate further, we analyzed various SAE features in depth and for each feature examined the top 100 samples with the strongest feature activations. We found that activation patterns of a substantial proportion of features can be interpreted in a meaningful way. Remarkably, these features were learned by SAE in an unsupervised manner.
While the previous features correspond to the entire genre classes in the dataset, some features capture categorizations that go beyond these predefined classes. For example, Figure 6(a) illustrates the activation pattern of a feature that activates on Japanese Anime items but remains largely inactive on Animation movies in other languages. The distinction is demonstrated in Figure 6(a) by highlighting the Animation genre in green and the Japanese original language in blue. When summing the activations of the top 100 samples, items with the Animation attribute account for $p_{top100}=0.633$ of all activations, while items labeled with Japanese as the original language comprise $p_{top100}=0.582$.
@@
There are multiple other curious representations learned by SAE. For instance, Figure 6(c) shows a feature that activates on non-English language movies. Other attributes of items, such as movie release year, were also reflected in SAE features. Figures 6(d) and 6(e) show two such features: the former activates on movies released around 2004-2005, while the latter captures movies from the 1980s, both in an apparently genre-agnostic manner.
@@ 8. Control
8.1. Examples
@@
For user 3634, there are a lot of Action and Thriller genres in the recommendations. After we set the activation of the Sci-Fi feature to +2, there are several movies from the original list and several new Sci-Fi movies. After we set the activation of the Sci-Fi feature to +5 (this is a very high value, exceeding most of the real examples in the dataset), all movies in the recommendation list belong to Sci-Fi. Interestingly, Action and Thriller remain among the most represented genres after the intervention. This means that the intervention performs a kind of soft control and shifts predictions toward Sci-Fi while taking into account other user interests. Thus this approach opens up an opportunity to mix the desired attributes to varying degrees.
8.2. Comprehensive Evaluation
In this section, we evaluate how modification of SAE activations influences genre presence in recommendations as well as recommendation metrics. We take top features, corresponding to Action, Children, Fantsy, Sci-Fi, and Western into our analysis.
@@ 8.2.2. Quality of recommendations
As one might expected, the intervention led to some decrease in the recommendation metrics. The patterns are similar for different features. The intervention with activation absolute values less than 2 or 3 is quite safe, as the decrease in NDCG@10 and Coverage@10 is less than 10% for most of the cases. Inter-list Diversity is also relatively stable in this range. Interventions with values greater than 5 should be performed with caution because metrics can decrease significantly. Interestingly, patterns for large negative values are similar to large positive values. When we suppress some genres, at the same time, we greatly increase the proportion of dissimilar genres, so the recommendations become less accurate and diverse.
@@ 9. Conclusion
In this work, we extended sparse autoencoders to the domain of sequential recommendations. We demonstrated that SAEs can learn meaningful and interpretable features, introducing and exploring novel interpretability metrics tailored to sequential recommendations. Additionally, we showed that these features can effectively control the model’s behavior, with their downstream effects aligning consistently with their interpretations. Furthermore, we proposed a flexible and adaptable approach for controlling sequential recommendation models, enabling previously unattainable modes of personalization and fine-tuning.
We have shown that methods of the rapidly developing field, Mechanistic Interpretability, can be extended to benefit the various domains that use transformer models. For future work in the recommendation domain, promising directions include exploring alternative approaches to SAE, such as JumpReLU (Rajamanoharan et al., 2024b) and TopK SAEs (Gao et al., 2024), and developing methods for controlling model behavior through inference-time interventions. Moreover, another methods of mechanistic interpretability, such as circuit discovery, causal scrubbing, etc.  (He et al., 2024) can also be adopted and extended intro recommendation domain.

### v2

@@ 7.2. Case Studies
At first, we examine the top features found in section 7.1.1. Figure 5 presents histograms of positive per-item feature activations for features exhibiting the highest correlation with Sci-Fi and Action. Histogram colors indicate whether the activation happened on an item of the associated genre. The proportion of such items increases notably with higher activation levels, which is reminiscent of the behavior of interpretable SAE features observed in NLP models (Cunningham et al., 2023; Bricken et al., 2023). We found the same pattern for all other genres. Activation of a given feature above some threshold typically indicates the presence of respective genre. However, the threshold varies between features, as well as the scale of activation values.
@@
To investigate further, we analyzed various SAE features in depth. For each feature we summed activations of the top 100 samples with the strongest responses and normalized by all activations in each sample, yielding the activation probability $p_{top100}$. We found that activation patterns of a substantial proportion of features can be interpreted in a meaningful way. Remarkably, these features were learned by SAE in an unsupervised manner.
Although the previous features correspond to the entire genre classes in the dataset, some features capture categorizations that go beyond these predefined classes. For example, Figure 6(a) illustrates the activation pattern of a feature that activates on Japanese Anime items but remains largely inactive on Animation movies in other languages. The distinction is demonstrated in Figure 6(a) by highlighting the Animation genre in green and the Japanese original language in blue. Items with the Animation attribute account for $p_{top100}=0.633$ of all activations, while items labeled with Japanese as the original language comprise $p_{top100}=0.582$.
@@
There are multiple other curious representations learned by SAE. For instance, Figure 6(c) shows a feature that activates on non-English language movies. Other attributes of items, such as movie release year, were also reflected in SAE features. Figures 6(d) and 6(e) show two such features: the former activates on movies released around 2004 – 2005, while the latter captures movies from the 1980s, both in an apparently genre-agnostic manner.
@@ 8. Control
In this section, we determine how the modification of SAE activation values influences the model predictions.
8.1. Examples
For user 3634, Action and Thriller genres predominate in the recommendations. After we set the activation of the Sci-Fi feature to +2, there are several movies from the original list and several new Sci-Fi movies. After we set the activation of the Sci-Fi feature to +5 (this is a very high value, exceeding most of the real examples in the dataset), all movies in the recommendation list belong to Sci-Fi. Interestingly, Action and Thriller remain among the most represented genres after the intervention. This means that the intervention performs a kind of soft control and shifts predictions toward Sci-Fi while taking into account other user interests. Thus this approach opens up an opportunity to mix the desired attributes to varying degrees.
@@ 8.1. Examples
8.2. Comprehensive Evaluation
In this section, we evaluate how modification of SAE activations influences genre presence in recommendations as well as recommendation metrics. We take top features, corresponding to Action, Children, Fantasy, Sci-Fi, and Western into our analysis.
@@ 8.2.2. Quality of recommendations
As one might expected, the intervention led to some decrease in the recommendation metrics. The patterns are similar for different features. The intervention with activation absolute values less than 2 is safe, as the decrease in NDCG@10 and Coverage@10 is less than 10% for most of the cases. Inter-list Diversity is also relatively stable in this range. Interventions with values greater than 5 should be performed with caution because metrics can decrease significantly. Interestingly, patterns for large negative values are similar to large positive values. When we suppress some genres, at the same time, we greatly increase the proportion of dissimilar genres, so the recommendations become less accurate and diverse.
The results of the intervention on the genre proportions and recommendation metrics for BERT4Rec and the Music4all dataset are provided in the appendix.
@@ 9. Conclusion
In this work, we extended sparse autoencoders to the domain of sequential recommendations. We demonstrated that SAEs can learn meaningful and interpretable features and introduced new interpretability metrics specifically tailored to sequential recommendation settings. Moreover, we showed that these learned features can be used to effectively and flexibly control model behavior, with their downstream effects aligning consistently with their interpretations.
Our findings highlight a new perspective on understanding and steering sequential recommender models. Beyond advancing interpretability, the proposed approach provides a practical mechanism for controllable personalization, enabling recommendation systems that are more transparent, adaptable, and user-aligned. Future work may explore alternative sparse autoencoder architectures, such as JumpReLU (Rajamanoharan et al., 2024b) and TopK SAEs (Gao et al., 2024), develop more refined control techniques, and investigate how interpretable latent representations can be integrated with broader recommendation objectives such as fairness, diversity, and long-term user satisfaction.
Appendix A Additional Results for BERT4Rec
For BERT4Rec, we perform the full set of experiments on the Movielens-20M dataset. First, we compute correlation, ROC AUC, and sensitivity between each SAE feature and genre and identify the top feature with the highest correlation, following the procedure from Section 7.1.1. Table 4 presents the metric values for these top features. Similar to GPTRec, the genres Horror, Animation, Sci-Fi, and Children are among those with the highest correlations. Overall, the absolute metric values are lower than for GPTRec. Figure 11 shows the correlations of these top features with all genres. As before, the features tend to be monosemantic, and the observed patterns align well with genre semantics.
Next, we perform the analysis from Section 7.1.3, comparing the interpretability of SAE features with that of the original transformer layer. Figure 12 displays the correlation values of the top transformer neurons and SAE features for each genre. Although the absolute correlations are lower than those for GPTRec, both for SAE and transformer neurons, the SAE features remain consistently more interpretable than the original transformer neurons.
Figure 11. Correlation between genres and top feature (the feature with maximum correlation) for each genre for BERT4Rec on the Movielens-20M dataset. One row corresponds to one feature and contains its correlations with all genres.
Figure 12. Comparison between the interpretability of SAE features and neurons of the original transformer layer for BERT4Rec on the Movielens-20M dataset. The maximum correlation for each genre is blue for the transformer layer and orange for SAE.
Table 4. Metrics for SAE features with maximum correlation for each genre for BERT4Rec on the Moivelens-20M dataset. The last row shows the averages across all genres, which are considered as overall interpretability metrics.
Genre Correlation ROC AUC Sensitivity Genre popularity
Horror 0.533 0.798 0.681 0.066
Animation 0.412 0.702 0.572 0.083
Sci-Fi 0.380 0.687 0.577 0.182
Children 0.367 0.630 0.543 0.095
Fantasy 0.326 0.621 0.497 0.135
IMAX 0.320 0.760 0.767 0.061
Documentary 0.293 0.751 0.721 0.011
Action 0.288 0.626 0.536 0.331
Crime 0.285 0.657 0.508 0.195
Musical 0.254 0.735 0.597 0.038
Drama 0.253 0.628 0.455 0.470
Film-Noir 0.247 0.775 0.747 0.011
Adventure 0.246 0.618 0.473 0.255
Comedy 0.246 0.600 0.436 0.311
War 0.221 0.628 0.488 0.061
Thriller 0.217 0.611 0.459 0.278
Romance 0.190 0.602 0.482 0.175
Mystery 0.179 0.625 0.535 0.099
Western 0.158 0.642 0.520 0.015
Mean 0.285 0.668 0.558 0.151
Finally, we conduct extensive experiments on model control, as described in Section 8.2. Figure 14 illustrates the effect of modifying a single SAE feature on the genre proportions and recommendation metrics. For Action, Sci-Fi, Fantasy, and Children, the genre proportions change as expected. Modifying the Western feature leads to a significant increase in the War genre, while Western itself is not strongly affected. This suggests that the feature is in fact associated with War, and that no meaningful feature was learned for Western. This is supported by the very low metric values for Western shown in Table 4.
Regarding the effect on recommendation quality, the trend is similar to GPTRec: interventions with activation values below 2 result in small changes to the metrics, while stronger interventions can lead to more noticeable degradation in performance.
(b) NDCG@10
(c) Coverage@10
(d) Inter-list Diversity@10
Figure 14. Effect of changing a SAE feature on the proportion of genres in recommendations (a) and the recommendation metrics(b-d) for BERT4Rec on the Movielens-20M dataset. For (a), each curve represents genre proportions for different activation values. Only genres with a large proportion of change are shown for clarity. For (b-d), each curve corresponds to one single feature that is responsible for a given genre.
Appendix B Additional Results for the Music4all Dataset
For the Music4All dataset, we perform the full set of experiments using GPTRec. Following the procedure from Section 7.1.1, we compute correlation, ROC AUC, and sensitivity between each SAE feature and genre and identify the top feature with the highest correlation. Table 5 presents the metric values for these top features. The feature corresponding to k-pop shows particularly high values across all metrics, indicating strong and specific alignment with this genre. While other features show lower values, many still reach levels comparable to those observed on the Movielens-20M dataset.
Figure 15 shows the correlations of these top features with all genres. The features appear even more monosemantic than in the movie domain. Several semantically meaningful patterns emerge: the rock feature correlates with alternative rock and indie rock, the rap feature with hip-hop, the synthpop feature with electropop, and the metal feature with both rock and alternative rock.
Table 5. Metrics for SAE features with maximum correlation for each genre for GPTRec on the Music4all dataset. The last row shows the averages across all genres, which are considered as overall interpretability metrics.
Genre Correlation ROC AUC Sensitivity Genre popularity
k-pop 0.883 0.983 0.971 0.033
classic rock 0.490 0.787 0.613 0.027
country 0.466 0.711 0.559 0.024
rap 0.457 0.723 0.546 0.070
hip hop 0.411 0.755 0.574 0.040
metal 0.342 0.815 0.719 0.022
indie rock 0.320 0.695 0.488 0.068
dream pop 0.317 0.735 0.531 0.022
rock 0.314 0.640 0.411 0.191
pop rock 0.288 0.619 0.381 0.033
experimental 0.268 0.656 0.398 0.025
alternative rock 0.258 0.673 0.422 0.069
pop 0.233 0.616 0.343 0.448
synthpop 0.230 0.635 0.402 0.046
electronic 0.227 0.596 0.314 0.107
soul 0.206 0.602 0.369 0.055
electropop 0.201 0.632 0.375 0.033
indie pop 0.162 0.584 0.308 0.057
folk 0.159 0.628 0.439 0.039
singer-songwriter 0.139 0.624 0.421 0.036
Mean 0.331 0.685 0.479 0.072
Figure 15. Correlation between genres and top feature (the feature with maximum correlation) for each genre for GPTRec on the Music4all dataset. One row corresponds to one feature and contains its correlations with all genres.
We also follow the procedure described in Section 8.2 and conduct extensive experiments on model control. Figure 17 illustrates how modifying a single SAE feature affects genre proportions and recommendation metrics. As with Movielens-20M, the proportion of the corresponding genre increases monotonically with activation value: it is near zero for large negative activations and reaches high levels for large positive activations. Additional patterns align with genre semantics and the correlations from Table 5. For example, increasing the rap feature also increases the proportion of hip-hop genre, increasing the metal feature increases the rock proportion, and increasing the rock feature boosts the proportion of alternative rock. In all these cases, the presence of pop genre tends to decrease.
The results regarding the effect on recommendation quality mirror previous findings: interventions with activation values below 2 lead to minimal changes in metrics, while stronger interventions can cause more significant degradation in performance.
(b) NDCG@10
(c) Coverage@10
(d) Inter-list Diversity@10
Figure 17. Effect of changing a SAE feature on the proportion of genres in recommendations (a) and the recommendation metrics(b-d) for GPTRec on the Music4all dataset. For (a), each curve represents genre proportions for different activation values. Only genres with a large proportion of change are shown for clarity. For (b-d), each curve corresponds to one single feature that is responsible for a given genre.
(a) Actor
(b) Company
(c) Country
(d) Director
Figure 18. Comparison between interpretability of SAE features and neurons of the original transformer layer for different item attributes. The maximum correlation for each movie attribute is blue for the transformer layer and orange for SAE.
Appendix C Analysis of Other Movie Attributes
To further explore the SAE feature space, we evaluate interpretability using additional item attributes beyond genres on the Movielens-20M dataset. Specifically, we consider the 20 most frequent actors, production companies, production countries, and directors. Following the procedure from Section 7.1.3, we compare the interpretability of SAE features with that of neurons from the original transformer layer.
Figure 18 shows the correlations between considered movie attributes and both the transformer neurons and SAE features. For all attributes, the association with SAE features is significantly stronger. Correlations are particularly high for actors and directors, moderate for production companies, and relatively low for most production countries.
Table 6 summarizes the results, reporting the minimum, mean, and maximum correlation values for both transformer neurons and SAE features. These observations indicate that the SAE captures diverse and meaningful patterns beyond genre-level information.
Table 6. Comparison of SAE features with original transformer layer neurons for different item attributes. Minimum, mean, and maximum correlation values for top features/neurons.
Attribute Transformer Layer SAE Features
min mean max min mean max
Actor 0.07 0.15 0.28 0.22 0.51 0.76
Company 0.07 0.13 0.25 0.15 0.35 0.72
Country 0.04 0.10 0.24 0.08 0.22 0.68
Director 0.06 0.13 0.28 0.22 0.52 0.79